OpenStudy (anonymous):

in triangle xyz the base is 11 and the area is 55m2. what is its height?

OpenStudy (anonymous):

area of a triangle = 1/2 * base * height plug in the numbers 55m^2 = 1/2 * 11 * height and solve for the height